OpenStudy (anonymous):

find the equation of the line that passes through the points (3,6) and (-5,9)

OpenStudy (anonymous):

y=mx+b to find m use \[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

m represents your slope

OpenStudy (anonymous):

Anyways, gtg game of thrones is on, gl!

OpenStudy (anonymous):

how is that the equation

OpenStudy (anonymous):

If you input your x and y point into that equation, it will give you the slope of the line you are trying to find.

OpenStudy (anonymous):

thats just the slope not equation

OpenStudy (luigi0210):

Or if you prefer, you could use the point-slope formula: \[\Large y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

so my points are 3,6 and -5,9

OpenStudy (anonymous):

Yes, now assign them x and y values. So x=3 y=6 and then do the same for the other set. Then you can input these into the point-slope formula to find the slope of the line you are trying to find.

OpenStudy (anonymous):

x1 is 3 and y1 is 6?

OpenStudy (anonymous):

Sure. Then let x=-5 and y=9. The important thing here is that you keep the sets of coordinates together. (x1,y1) and (x,y). Now try putting these numbers into the point-slope equation that was posted above.

OpenStudy (anonymous):

ok here is what i have 9-6 =m(-5-3)

OpenStudy (anonymous):

Excellent. Now solve for m (this is giving you the slope).

OpenStudy (anonymous):

-0.375

OpenStudy (anonymous):

So what we have now is the slope of our line (m=-0.375 or -3/8) and we know that the line intersects the points (3,6) and (-5,9). We actually only need to know 1 set of coordinates. So pick the set and enter that, and the slope, into the point-slope form.

OpenStudy (anonymous):

how i dont get it

OpenStudy (anonymous):

The number you just solved for (-0.375) is the slope of the line. It tells you the rise (the change in y values) divided by the run (the change in x values).

OpenStudy (anonymous):

y-3=0.375(x-6)

OpenStudy (anonymous):

Yes. Now you could leave the equation like that, or you might want to solve for y so that it is in slope-intercept form (y=mx+b).

OpenStudy (anonymous):

Does your question specify how they want you to answer the question?

OpenStudy (anonymous):

all it says is this: Find the equation of the line that passes through the points (3,6) and (-5,9)

OpenStudy (anonymous):

They probably want you to write it in y=mx+b form (at least that is what I used to do). So solve your equation for y and you should be done! :)

OpenStudy (anonymous):

how would i solve for y?

OpenStudy (anonymous):

By solving for y I just mean do some algebra and move the equation so that y is isolated on the left hand side.

OpenStudy (anonymous):

this is confusing for me sorry i probably sound like an idiot haha

OpenStudy (anonymous):

No you don't! This stuff can be confusing. Just try to understand each step as opposed to just plugging in numbers into random formulas and you should be fine!

OpenStudy (anonymous):

so move y to the other side i would take y and put it like this -3= 0.375(x-6)-y

OpenStudy (anonymous):

I think we made a mistake

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay when you put your coordinates and slope into the point-slope equation, you mixed up the x and y values. So instead of y-3=-0.375(x-6) it should be y-6=-0.375(x-3)

OpenStudy (anonymous):

Because the equation is y-y1=m(x-x1)

OpenStudy (anonymous):

ok so -6=0.375(x-3)-y

OpenStudy (anonymous):

Yup but you do not need to move the y over to the right hand side. We want to isolate y (which means it will say y=.....)

OpenStudy (anonymous):

So try adding the 6 to both sides instead

OpenStudy (anonymous):

what do i do with the (x-3)

OpenStudy (anonymous):

Try multiplying it by the -0.375

OpenStudy (anonymous):

6.375(x-3)

OpenStudy (anonymous):

Ah, there is a bit of an algebra mistake there

OpenStudy (anonymous):

y-3=-0.375(x-6) add 3 to both sides y=-0.375(x-6)+3

OpenStudy (anonymous):

You cannot add the -0.375 and the 3 together directly because the -0.375 is being multiplied by the (x-6)

OpenStudy (anonymous):

so would it be -6+3?

OpenStudy (anonymous):

Nope! the -6 is in the brackets, you can simply add the +3 to the end of the equation.

OpenStudy (anonymous):

so that is the whole equation

OpenStudy (anonymous):

y=-0.375(x-6)+3

OpenStudy (anonymous):

Yes, but to answer in slope-intercept form (y=mx+b) you should multiply the -0.375 by the (x-6)

OpenStudy (anonymous):

1.875

OpenStudy (anonymous):

I'm not sure what you did there.

OpenStudy (anonymous):

ok so how do i multiply it? since i cant multiply x would i just multiply the 6

OpenStudy (anonymous):

6 and -0.375

OpenStudy (anonymous):

You can multiply it by the x. You would multiply it by the x and by the -6

OpenStudy (anonymous):

i got 1.875

OpenStudy (anonymous):

What did you do to get that answer?

OpenStudy (anonymous):

-0.375 x 1 x 6

OpenStudy (anonymous):